OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192. Select one: A. 134,217,728 B. 33,554,432 C. -33,554,432 D. -134,217,728

zepdrix (zepdrix):

Umm ok so here's what I'm thinking...

zepdrix (zepdrix):

To get from one term to the next, we multiply by some `common ratio` yes? Let's call it r. So to get from a1 to a2,\[\Large\rm a_2=r a_1\]We multiply by r, yah? To get from a2 to a3, we multiply by r again,\[\Large\rm a_3=r a_2\]\[\Large\rm a_3=r^2a_1\]So we if keep going along like this, we eventually get,\[\Large\rm a_6=r^5a_1\]And we can plug in the information they gave us to try and find our common ratio:\[\Large\rm -8192=8r^5\]Solving for r:\[\Large\rm r=\sqrt[5]{\frac{-8192}{8}}\]And then to get the 12th term of the sequence, we just take our first term and multiply it by the common ratio 11 times. \[\Large\rm a_{12}=a_1 \left(r\right)^{11}\]

OpenStudy (anonymous):

(8)^11

zepdrix (zepdrix):

No we're raising the `common ratio (r)` to the 11th power. You have to solve for r using the information I gave you above. Then multiply that huge number by 8.

OpenStudy (anonymous):

Is the answer C?

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

Sweet Thanks!